# Predictable Scale: Part I, Step Law – Optimal Hyperparameter Scaling Law in Large Language Model Pre-training

### v6

@@ 1 Introduction
To avoid exhaustive searches, the community has advanced hyperparameter-transfer rules. $\mu$P Yang2022 first established learning rate scaling with model width, later extended to depth and other architectures Everett2024 ; Lingle2024 ; Blake2024 . Empirical scaling laws then linked learning rate and batch size to model size kaplan_scaling_2020 , refined for dense models in subsequent work Bjorck2024 ; DeepSeek-AI2024 ; Porian2024 ; Hu2024 . Recently, studies have explored these interactions in Mixture-of-Experts (MoE) models Du2021 ; Fedus2021 ; Wang2024 ; Ludziejewski2025 .
@@ 2 Related Works
[…]
@@ 3.4.1 Scaling Laws
Figure 5: Empirical vs. predicted HP scaling. (a) Scatter points indicate empirical optimal learning rate vs. batch size for model scale $N$; (b) Analogous results for dataset scale D. Curves show our hp-scaling law predictions, with shaded regions representing parameter uncertainty bounds from the sampling-based fitting strategy. Both plots use double logarithmic scaling (1,912 training samples).
@@ 3.4.3 Comparisons with Existing Approaches
With the fitted scaling laws, we directly extrapolate to the test point ($N=1$B, $D=100$B) for comparison. As shown in Fig. 1 and detailed in Appendix A.7, our method consistently finds solutions near the global optimum. In contrast, other approaches show significant deviation.This gap mainly comes from their modeling limitations.
@@
We further test on a 6.5B-parameter MoE model with two dataset sizes: $D=1.0\times 10^{10}$ and $1.3\times 10^{11}$ (see configurations 17 and 18 in Appendix A.3). For each setting, we use Porian Law, DeepSeek Law, and our Step Law to predict LR and BS from $(N,D)$, and train accordingly. We exclude $\mu$Yang2022 as it only predicts LR and does not handle BS.Table 3 shows that our Step Law consistently achieves the lowest validation smooth loss. This confirms that our method generalizes well to large, sparse MoE architectures and remains robust under diverse training regimes.
@@ A.1 Notation
  • $N$: Number of non-embedding parameters in the model.
  • $\hat{N}$: Total number of parameters in the model, which excludes embedding layer but includes the models head layer.
@@ A.2 Model Scale Dominates Optimal Hyperparameter Selection Over Computational Complexity
$d_{{model}}$ $d_{{ff}}$ $N_{{head}}$ $N_{{layer}}$ $\eta(N,D)$ $B(N,D)$ $D$ $N$ $M$
@@
1280 12264 10 8 $1.95\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.83\times 10^{9}$
1280 6280 10 14 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.02\times 10^{9}$
1536 9600 12 8 $9.77\times 10^{-4}$ $131{,}072$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.88\times 10^{9}$
1536 7264 12 10 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.95\times 10^{9}$
1536 4608 12 14 $9.77\times 10^{-4}$ $131{,}072$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.10\times 10^{9}$
2048 6000 16 8 $9.77\times 10^{-4}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.98\times 10^{9}$
2048 4256 16 10 $9.77\times 10^{-4}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.08\times 10^{9}$
2048 2256 16 14 $9.77\times 10^{-4}$ $262{,}144$ $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.28\times 10^{9}$
@@
1280 12608 10 8 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.40\times 10^{8}$ $2.89\times 10^{9}$
1280 5888 10 14 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.08\times 10^{8}$ $2.89\times 10^{9}$
1536 9656 12 8 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.31\times 10^{8}$ $2.89\times 10^{9}$
1536 7040 12 10 $1.38\times 10^{-3}$ $262{,}144$ $8.00\times 10^{9}$ $4.19\times 10^{8}$ $2.89\times 10^{9}$
1536 4056 12 14 $9.77\times 10^{-4}$ $262{,}144$ $8.00\times 10^{9}$ $3.94\times 10^{8}$ $2.89\times 10^{9}$
2048 5704 16 8 $9.77\times 10^{-4}$ $262{,}144$ $8.00\times 10^{9}$ $4.15\times 10^{8}$ $2.89\times 10^{9}$
2048 3744 16 10 $6.91\times 10^{-4}$ $131{,}072$ $8.00\times 10^{9}$ $3.98\times 10^{8}$ $2.89\times 10^{9}$
2048 1504 16 14 $6.91\times 10^{-4}$ $131{,}072$ $8.00\times 10^{9}$ $3.64\times 10^{8}$ $2.89\times 10^{9}$
Table 4: Model configurations and results for constant $N$ and constant $M$ experiments. The first group (top) maintains constant parameter count $N$approx $4.29\times 10^{8}$, while the second group (bottom) maintains constant computational complexity $M$approx $2.89\times 10^{9}$. $M$: non-embedding FLOPs/token.
@@
[…]
@@ A.3 Model Structural Parameters
Model $N$ $D$ $d_{{model}}$ $d_{{ff}}$ $N_{{head}}$ $N_{{layer}}$
@@
Model $N$ $D$ $d_{{model}}$ $N_{{head}}$ $N_{{layer}}$ $N_{{expert}}$ $d_{{moe}}$ Top-$k$ $N_{a}$

### v7

@@ 1 Introduction
[…]
@@ 2 Related Works
Theory-driven methods began with $\mu\text{P}$’s learning rate rules for varying model widths Yang2022 . Subsequent work extended these rules to model depth and other architectural variants Everett2024 ; Lingle2024 ; Blake2024 ; Bordelon2023 . All of these approaches require custom initialization or modified attention mechanisms. They do not address variations in data recipe, sparsity level or dataset size, nor do they predict batch size.
@@ 3.4.1 Scaling Laws
Figure 5: Empirical vs. predicted HP scaling. (a) Scatter points indicate empirical optimal learning rate vs. batch size for model scale NN; (b) Analogous results for dataset scale D. Curves show our hp-scaling law predictions, with shaded regions representing parameter uncertainty bounds from the sampling-based fitting strategy. Both plots use double logarithmic scaling (1,912 training samples).
@@ 3.4.3 Comparisons with Existing Approaches
With the fitted scaling laws, we directly extrapolate to the test point (N=1N=1B, D=100D=100B) for comparison. As shown in Fig. 1 and detailed in Appendix A.7, our method consistently finds solutions near the global optimum. In contrast, other approaches show significant deviation.This gap mainly comes from their modeling limitations.
@@
[…]
@@ A.1 Notation
  • NN: Number of non-embedding parameters in the model.
  • $\hat{N}$: Total number of parameters in the model, which excludes embedding layer but includes the model’s head layer.
@@ A.2 Model Scale Dominates Optimal Hyperparameter Selection Over Computational Complexity
$d_{{model}}$ $d_{{ff}}$ $N_{{head}}$ $N_{{layer}}$ $\eta(N,D)$ $B(N,D)$ $D$ NN $M$
@@
1280 12264 10 8 $1.95\times 10^{-3}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.83\times 10^{9}$
1280 6280 10 14 $1.38\times 10^{-3}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.02\times 10^{9}$
1536 9600 12 8 $9.77\times 10^{-4}$ 131,072131{,}072 $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.88\times 10^{9}$
1536 7264 12 10 $1.38\times 10^{-3}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.95\times 10^{9}$
1536 4608 12 14 $9.77\times 10^{-4}$ 131,072131{,}072 $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.10\times 10^{9}$
2048 6000 16 8 $9.77\times 10^{-4}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.29\times 10^{8}$ $2.98\times 10^{9}$
2048 4256 16 10 $9.77\times 10^{-4}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.08\times 10^{9}$
2048 2256 16 14 $9.77\times 10^{-4}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.29\times 10^{8}$ $3.28\times 10^{9}$
@@
1280 12608 10 8 $1.38\times 10^{-3}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.40\times 10^{8}$ $2.89\times 10^{9}$
1280 5888 10 14 $1.38\times 10^{-3}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.08\times 10^{8}$ $2.89\times 10^{9}$
1536 9656 12 8 $1.38\times 10^{-3}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.31\times 10^{8}$ $2.89\times 10^{9}$
1536 7040 12 10 $1.38\times 10^{-3}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.19\times 10^{8}$ $2.89\times 10^{9}$
1536 4056 12 14 $9.77\times 10^{-4}$ 262,144262{,}144 $8.00\times 10^{9}$ $3.94\times 10^{8}$ $2.89\times 10^{9}$
2048 5704 16 8 $9.77\times 10^{-4}$ 262,144262{,}144 $8.00\times 10^{9}$ $4.15\times 10^{8}$ $2.89\times 10^{9}$
2048 3744 16 10 $6.91\times 10^{-4}$ 131,072131{,}072 $8.00\times 10^{9}$ $3.98\times 10^{8}$ $2.89\times 10^{9}$
2048 1504 16 14 $6.91\times 10^{-4}$ 131,072131{,}072 $8.00\times 10^{9}$ $3.64\times 10^{8}$ $2.89\times 10^{9}$
Table 4: Model configurations and results for constant NN and constant $M$ experiments. The first group (top) maintains constant parameter count NN $\approx$ $4.29\times 10^{8}$, while the second group (bottom) maintains constant computational complexity $M$ $\approx$ $2.89\times 10^{9}$. $M$: non-embedding FLOPs/token.
@@
The experimental results reveal several key findings: (i) Models with constant $N$ demonstrate remarkably consistent optimal hyperparameter regions, with minimal variation in minimum loss values (ranging from 2.4294 to 2.4776) despite significant architectural differences. (ii) The constant M experiments show slightly more variation in optimal hyperparameter regions and minimum loss values (ranging from 2.4346 to 2.5089), suggesting that parameter count $N$ may be a more robust indicator for hyperparameter selection than computational complexity $M$. (iii) Across both experimental groups, the optimal learning rates typically fall within a narrow range ($6.91\times 10^{-4}$ to $1.95\times{-3}$), and batch sizes cluster around either 131,072 or 262,144, regardless of the specific architectural configuration.
@@ A.3 Model Structural Parameters
Model NN $D$ $d_{{model}}$ $d_{{ff}}$ $N_{{head}}$ $N_{{layer}}$
@@
Model NN $D$ $d_{{model}}$ $N_{{head}}$ $N_{{layer}}$ $N_{{expert}}$ $d_{{moe}}$ Top-$k$ $N_{a}$